OpenStudy (anonymous):

How can you make x the subject in this --> y=x^3-2x^2

OpenStudy (amistre64):

declare a state of insolvency and elect a new parliment ... is just my first instinct

OpenStudy (mr.math):

LOL!

OpenStudy (mr.math):

New president?

jhonyy9 (jhonyy9):

y=x2(x-2) - hope you will like it

OpenStudy (mr.math):

It's very difficult to solve for x.

OpenStudy (amistre64):

can we make it a perfect cube?

OpenStudy (mr.math):

We can't!

OpenStudy (anonymous):

I knew there was something wrong!! But how can you then solve this question..if rate of change of y is 4 units per second, find rate of change at x at the given instant y=x^3-2x^2, x=3

OpenStudy (amistre64):

rate of chage is derivatives

OpenStudy (amistre64):

y' = 3x^2 -4x gives the rate of change at any given moment

OpenStudy (mr.math):

So y and x are both functions in t?

OpenStudy (anonymous):

y'= 3 x^2 x' -4 x x'

OpenStudy (amistre64):

if x(t) = t; then x' = t' = 1

OpenStudy (mr.math):

If they are both functions of t, then do as Imran just did above, and solve it for x'.

OpenStudy (anonymous):

y'=4 x=3

OpenStudy (anonymous):

but am i not supposed to use the chain rule?

OpenStudy (anonymous):

we did

OpenStudy (anonymous):

and is the final answer the rate of change of x?